OpenStudy (anonymous):

3 1/2 - 2/3 need to see the work

OpenStudy (anonymous):

3 1/2 - 2/3

OpenStudy (anonymous):

first turn the 3 1/2 into improper

OpenStudy (anonymous):

do you know how do do that?

OpenStudy (anonymous):

hello

hero (hero):

 Alternatively \[3 + \frac{1}{2} - \frac{4}{6} \\= 3 + \frac{3}{6} - \frac{4}{6} \\ = 3 - \frac{1}{6} \\= 2 + 1 - \frac{1}{6} \\= 2 + \frac{6}{6} - \frac{1}{6} \\= 2 + \frac{5}{6}\]